# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v2

@@ II Magnetic colloidal assemblies contain prominent edge flows
Magnetically assembled sheets and clusters are created by subjecting a suspension of 1 $\mu$m super-paramagnetic particles (Dynabeads MyOne Carboxylic Acid, Invitrogen) in 10mM NaCl to a continuously rotating magnetic field. The suspension is held within an epoxy sealed chamber made of negatively charged glass slides and a Parafilm spacer. Since the particles are denser than the fluid and also negatively charged, they settle to a quasi-two-dimensional plane just above the substrate. The magnetic field is created by running a sine and cosine wave through two pairs of magnetic coils set orthogonal to one another, see Fig. 1(a). Upon application of a 20Hz field, the particles acquire magnetic dipoles and begin to interact according to the time-averaged interaction potential for super-paramagnetic particles in a continuously rotating magnetic field, that was theoretically derived and experimentally measured in [16]. This two-dimensional interaction potential, shown on the right of Fig. 1(a) for different field strengths, is composed of an electrostatic short-range repulsion and a long-range magnetic attraction. The short-range repulsion can be tuned by the salt solution, while the long-range attraction is tuned by the magnetic field strength, $H_{0}$. Under this relatively fast frequency field, particle assemblies with well defined, tunable order are possible [45].
@@
At the edge of voids and on the perimeter of clusters, we observe that particles exhibit edge flows, as shown by the particle trajectories in Figs. 1(b) and (c). Edge flows are composed of the particles along the edge which translate faster than particles in the bulk, as reflected in the mean speed, $v_{\textnormal{mean}}$, computed in the IMARIS software as the total path traveled over the length of time tracked. Note that all particles have a constant background thermal speed computed as an average path traveled due to the Brownian motion. The faster translation at the edge occurs regardless of magnetic field strength. Along voids, all edge particles translate but only a few particles move significantly at any one point in time. Meanwhile, the polycrystalline bulk remains fairly stationary, beyond thermal motion, over the course of an experiment (approximately 1 minute). In contrast, all particles along the edge of clusters move continuously and the entire cluster rotates as a result.
@@ III Topological origin of edge flows
where the length scale $\delta=\sqrt{(\eta+\eta_{R})/\Gamma_{v}}$ is determined by the two viscosities and friction with the substrate, $\Gamma_{v}$ (see SM for details). Notably, this operator has the same form as the square of a well-known topological Hamiltonian, the Dirac Hamiltonian (see SM for details). Eigenvectors of this Hamiltonian have a winding number known as the Chern number, which yields edge states robust to defects and random perturbations [23, 50].
At the interface between particle-rich and particle-poor regions, the Chern number changes its value, inducing a state with zero eigenvalue localized at the interface [32]. This state is exactly the steady state solution of Eq. (1), since these operators have the same zero eigenvalue solution. This solution yields a topologically-protected steady state where vorticity $\zeta(x)$ decays exponentially away from the edge, $x$. To obtain the particles speed $v(x)$ along the edge, we integrate $\zeta(x)$ (see SM for details) to obtain,
@@
Figure 2(b) shows Eq. 2 fit to the experimentally measured $v_{\mathrm{mean}}(x)$ for three voids under increasing values of magnetic field strength. At all field strengths, the average thermal motion is significant, corresponding to the value at which the edge speeds plateau ($v_{\mathrm{th}}$). However, despite this significant thermal motion, we find the R-squared statistic of fits to be 0.86, 0.89, and 0.83 for the field strengths $H_{0}=1.0$, $1.1$, and $1.2$ mT, respectively, showing that the topological edge flows are robust to strong noise. As the magnetic field increases, the magnetic interactions are expected to dominate over the hydrodynamic effects. This magnetic domination is confirmed by the decrease in the fitted maximal edge speed with field strength, showcased in Fig. 2(c) for five ($1.0$mT) , seven ($1.1$mT), and six voids ($1.2$mT). Since the increased magnetic interactions promote a stronger attraction between particles that densifies the cluster, the effect of hydrodynamics is reduced, thereby reducing the strength of edge flows. This is consistent with the decrease in $v_{\mathrm{th}}$ with the magnetic field strength in Fig. 2(b), since stronger magnetic interactions also decrease random motion.
@@ IV Edge flows induce cluster rotation
[…]
@@
To evaluate the importance of the rotational component in Eq. 4, we compare the fit with (Eq. 4) and without a rotational term (Eq. (2)) by looking at the fitted maximal edge speed with field strength. As shown in Fig. 3(c), five ($1.0$mT), nine ($1.1$mT)), and seven ($1.2$mT)) clusters are analyzed for each case. When the fit with the rotational term is considered, the maximal edge speeds of the clusters decrease with field strength (red line in Fig. 3(c)), in line with the trend of the voids. In contrast, when the rotational term is not considered, the opposite trend occurs and the maximal edge speed increases with field strength (purple line in Fig. 3(c)). Since the balance of hydrodynamic and magnetic interactions should remain constant with magnetic field, independent of geometry, we find the rotational term is necessary to properly describe edge flows in the cluster geometry. This indicates significant shear induced rotation of clusters in contrast to voids that are fixed in place and cannot rotate.
@@ VII Acknowledgments
Supplemental Material to: Topological edge flows drive macroscopic re-organization in magnetic colloids
Aleksandra Nelson,1,* Dana M. Lobmeyer,2,* Sibani L. Biswal,2 and Evelyn Tang1,3
1Center for Theoretical Biological Physics, Rice University, Houston TX, USA
2
Department of Chemical and Biomolecular Engineering, Rice University, Houston TX, USA
3
Department of Physics and Astronomy, Rice University, Houston TX, USA
These authors contributed equally to this work
(Dated: January 24, 2025)
A Hydrodynamic theory of colloidal assemblies
@@
$\rho\partial_{t}v_{i}$ $=-\partial_{i}p+\eta\nabla^{2}v_{i}+\eta_{R}\epsilon_{ij}\partial_{j}(2\omega_{p}-\zeta)-\Gamma_{v}v_{i},$ (S1)
$\mathcal{I}\partial_{t}\omega_{p}$ $=D_{\omega}\nabla^{2}\omega_{p}-2\eta_{R}(2\omega_{p}-\zeta)-\Gamma_{\omega}\omega_{p}+T,$ (S2)
@@
$\eta\nabla^{2}\zeta-\eta_{R}\nabla^{2}(2\omega_{p}-\zeta)-\Gamma_{v}\zeta=0,$ (S3)
$D_{\omega}\nabla^{2}\omega_{p}-2\eta_{R}(2\omega_{p}-\zeta)-\Gamma_{\omega}\omega_{p}+T=0,$ (S4)
where Eq. (S3) is obtained by taking a curl of Eq. (S1). In the bulk of the colloidal assembly the solution is
$\bar{\zeta}$ $=0,$ (S5)
$\bar{\omega}_{p}$ $=\frac{T}{\Gamma_{\omega}+4\eta_{R}},$ (S6)
@@
Near the boundary, the angular velocity and vorticity fields may vary to fulfill the boundary conditions. To simplify calculations, we assume that the particles spin much faster than they rotate around each other, e.g. $\omega_{p}\gg\zeta$. As we will see later, this assumption is reasonable for assemblies of magnetic colloids. Then the angular momentum remains a constant given by Eq. (S6), while the vorticity is given by the equation:
$\nabla^{2}\zeta-\delta^{-2}\zeta=0,\quad\delta=\sqrt{\frac{\eta+\eta_{R}}{\Gamma_{v}}}.$ (S7)
B Topological theory and edge solution
Notably, the operator in Eq. (S7) has the same form as the square of a well-known topological Hamiltonian, the Dirac Hamiltonian, i.e., $(-\nabla^{2}+\delta^{-2})I_{2}=\mathcal{H}_{d}^{2}$, where $I_{2}$ is a rank two identity matrix [32]. The Dirac Hamiltonian is $\mathcal{H}_{d}=m\sigma_{1}-i\partial_{x}\sigma_{2}-i\partial_{y}\sigma_{3}$, where $\sigma_{1,2,3}$ are the Pauli matrices and $m=\pm\delta^{-1}$. The eigenvectors of $\mathcal{H}_{d}$ have a winding number known as the Chern number, which yields edge states robust to defects and random perturbations [23, 50].
At the interface between particle-rich and particle-poor regions, the Chern number changes its value, inducing a state with zero eigenvalue localized at the interface [32]. This state is exactly the steady state solution of Eq. (S7), since the two operators $\mathcal{H}_{d}$ and $\mathcal{H}_{d}^{2}$ have the same zero eigenvalue solution. This solution yields a topologically-protected steady state where vorticity $\zeta(x)$ decays exponentially away from the edge, $x$.
Figure S5: (a) Colloidal cluster forming a semi-infinite plane. (b) Colloidal cluster with a generic boundary. At each point of the boundary define its curvature and a polar coordinate system.
To analytically find this exponentially decaying solution of Eq. (S7), we adopt another simplification, namely that the boundary of colloidal assembly is close to being flat. Then we can consider the colloidal assembly as a semi-infinite plane at $x>0$ with $x=0$ line being its boundary (see Fig. S5(a)). Then, the solution is $y$-independent and is
$\zeta(x)=\zeta_{0}e^{-x/\delta}.$ (S8)
@@ B Topological theory and edge solution
$\sigma_{yy}\big{|}_{x=0}$ $=-p=0,$ (S9)
$\sigma_{xy}\big{|}_{x=0}$ $=2\omega_{p}\eta_{R}+(\eta+\eta_{R})\zeta_{0},$ (S10)
@@
$\zeta_{0}=\frac{2\eta_{R}\omega_{p}}{\eta+\eta_{R}}.$ (S11)
@@
[…]
@@
Figure S6: Edge flow profile modeled in a cluster of radius $R$ as a function of the radial coordinate $r$. Red line shows an exponential decay with decay length $\delta$. Blue dashed line shows a modified Bessel function of the first kind with decay length $\delta$. Both functions are normalized to their values at the edge of the cluster. The ratio $R/\delta=5$.
C Solution at curved boundary
In this section we find the edge flow solution in a more general case of a curved boundary with a curvature $\kappa=1/R$. We introduce polar coordinates $r,\varphi$ with the center of coordinate located at the center of the circle with radius $R$ [cf. Fig. S5(b)]. We redefine linear velocity in these coordinates as $v_{r}=\dot{r}$ and $v_{\varphi}=r\dot{\varphi}$ and the vorticity is expressed as $\zeta=1/r\,v_{\varphi}+\partial_{r}v_{\varphi}-1/r\partial_{\varphi}v_{r}$. Rewriting Eq. (S7) in polar coordinates we get
$r^{2}\partial_{r}^{2}v_{\varphi}+r\partial_{r}v_{\varphi}-(1+r^{2}/\delta^{2})v_{\varphi}=0,$ (S13)
@@ C Solution at curved boundary
$v_{\varphi}(r)=\begin{cases}AI_{1}(r/\delta),&\textrm{convex boundary}\\ BK_{1}(r/\delta),&\textrm{concave boundary}\end{cases}$ (S14)
@@
$\sigma_{rr}\big{|}_{r=R}$ $=-p=\gamma/R,$ (S15)
$\sigma_{\varphi r}\big{|}_{r=R}$ $=(\eta_{R}+\eta)\partial_{r}v_{\varphi}+(\eta_{R}-\eta)/r\,v_{\varphi}-2\eta_{R}\omega_{p}\big{|}_{r=R}=0,$ (S16)
@@
$v_{\varphi}(r)=\begin{cases}\frac{\eta_{R}I_{1}(r/\delta)}{\eta I_{2}(R/\delta)+\eta_{R}I_{0}(R/\delta)}2\omega_{p}\delta,&\textrm{convex boundary}\\ -\frac{\eta_{R}K_{1}(r/\delta)}{\eta K_{2}(R/\delta)+\eta_{R}K_{0}(R/\delta)}2\omega_{p}\delta,&\textrm{concave boundary}\end{cases}$ (S17)
Note, that the modified Bessel function also forms an exponential decay. If the cluster radius is a few times larger than the decay length of the edge flow, the Bessel function is very close to the exponential decay, as we show in Fig. S6. All clusters and voids in our experiments have a radius to decay length ratio $R/\delta\gtrsim 5$, meaning they are large enough to be modeled by the exponential function in Eq. (2) in the main text.
D Derivation of bulk cluster rotation
In this section we derive the bulk angular velocity of a cluster induced by shear stress from the edge flows. We assume that only a thin outer layer of particles moves with velocity $v_{\varphi}(R)$. The friction between this layer and the bulk of the cluster imposes a torque
$\tau_{1}=2\pi R\cdot v_{\varphi}(R)\eta\cdot R,$ (S18)

### v3

@@ II Magnetic colloidal assemblies contain prominent edge flows
[…]
@@
[…]
In addition, all particles have a constant background thermal speed computed as an average path traveled due to the Brownian motion. Brownian particles display structure on all length scales, so while the total path of a Brownian trajectory depends on the sampling rate, this rate will affect the magnitude of the plateau but not the existence of the plateau itself. We estimate our noise plateaus to be about 2 $\mu\mathrm{m}/\mathrm{s}$, which is adequately captured by our sampling rate as discussed above. Notably, most of the data lie in the exponentially-decaying portion, from which we can see that the faster translation at the edge occurs regardless of magnetic field strength. Along voids, all edge particles translate but only a few particles move significantly at any one point in time. Meanwhile, the polycrystalline bulk remains fairly stationary, beyond thermal motion, over the course of an experiment (approximately 1 minute). In contrast, all particles along the edge of clusters move continuously and the entire cluster rotates as a result.
@@ III Topological origin of edge flows
where the length scale $\delta=\sqrt{(\eta+\eta_{R})/\Gamma_{v}}$ is determined by the two viscosities and friction with the substrate, $\Gamma_{v}$ (see Appendix A for details). Notably, this operator has the same form as the square of a well-known topological Hamiltonian, the Dirac Hamiltonian (see Appendix B for details). Eigenvectors of this Hamiltonian have a winding number known as the Chern number, which yields edge states robust to defects and random perturbations [23, 50].
At the interface between particle-rich and particle-poor regions, the Chern number changes its value, inducing a state with zero eigenvalue localized at the interface [32]. This state is exactly the steady state solution of Eq. (1), since these operators have the same zero eigenvalue solution. This solution yields a topologically-protected steady state where vorticity $\zeta(x)$ decays exponentially away from the edge, $x$. To obtain the particles speed $v(x)$ along the edge, we integrate $\zeta(x)$ (see Appendix B for details) to obtain,
@@
Figure 2(b) shows Eq. 2 fit to the experimentally measured $v_{\mathrm{mean}}(x)$ for three voids under increasing values of magnetic field strength. At all field strengths, the average thermal motion is significant, corresponding to the value at which the edge speeds plateau ($v_{\mathrm{th}}$). However, despite this significant thermal motion, we find the R-squared statistic of fits to be 0.86, 0.89, and 0.83 for the field strengths $H_{0}=1.0$, $1.1$, and $1.2$ mT, respectively, showing that the topological edge flows are robust to strong noise. As the magnetic field increases, the magnetic interactions are expected to dominate over the hydrodynamic effects. This magnetic domination is confirmed by the decrease in the fitted maximal edge speed with field strength, showcased in Fig. 2(c) for five (1.01.0mT) , seven (1.11.1mT), and six voids (1.21.2mT). Since the increased magnetic interactions promote a stronger attraction between particles that densifies the cluster, the effect of hydrodynamics is reduced, thereby reducing the strength of edge flows. This is consistent with the decrease in $v_{\mathrm{th}}$ with the magnetic field strength in Fig. 2(b), since stronger magnetic interactions also decrease random motion.
@@ IV Edge flows induce cluster rotation
where $v_{i}$ is the speed of the particle of interest, $v_{\mathrm{interior}}$ is the speed of the neighboring particle most interior, $x$ is the distance between the two particles and $\langle\rangle$ is a time average. Since the viscosity is assumed constant, it does not appear in the expression above.
@@
To evaluate the importance of the rotational component in Eq. 4, we compare the fit with (Eq. 4) and without a rotational term (Eq. (2)) by looking at the fitted maximal edge speed with field strength. As shown in Fig. 3(c), five (1.01.0mT), nine (1.11.1mT)), and seven (1.21.2mT)) clusters are analyzed for each case. When the fit with the rotational term is considered, the maximal edge speeds of the clusters decrease with field strength (red line in Fig. 3(c)), in line with the trend of the voids. In contrast, when the rotational term is not considered, the opposite trend occurs and the maximal edge speed increases with field strength (purple line in Fig. 3(c)). Since the balance of hydrodynamic and magnetic interactions should remain constant with magnetic field, independent of geometry, we find the rotational term is necessary to properly describe edge flows in the cluster geometry. This indicates significant shear induced rotation of clusters in contrast to voids that are fixed in place and cannot rotate.
@@ VII Acknowledgments
Appendix A Hydrodynamic theory of colloidal assemblies
@@
$\rho\partial_{t}v_{i}$ $=-\partial_{i}p+\eta\nabla^{2}v_{i}+\eta_{R}\epsilon_{ij}\partial_{j}(2\omega_{p}-\zeta)-\Gamma_{v}v_{i},$ (5)
$\mathcal{I}\partial_{t}\omega_{p}$ $=D_{\omega}\nabla^{2}\omega_{p}-2\eta_{R}(2\omega_{p}-\zeta)-\Gamma_{\omega}\omega_{p}+T,$ (6)
@@
$\eta\nabla^{2}\zeta-\eta_{R}\nabla^{2}(2\omega_{p}-\zeta)-\Gamma_{v}\zeta=0,$ (7)
$D_{\omega}\nabla^{2}\omega_{p}-2\eta_{R}(2\omega_{p}-\zeta)-\Gamma_{\omega}\omega_{p}+T=0,$ (8)
where Eq. (7) is obtained by taking a curl of Eq. (5). In the bulk of the colloidal assembly the solution is
$\bar{\zeta}$ $=0,$ (9)
$\bar{\omega}_{p}$ $=\frac{T}{\Gamma_{\omega}+4\eta_{R}},$ (10)
@@
Near the boundary, the angular velocity and vorticity fields may vary to fulfill the boundary conditions. To simplify calculations, we assume that the particles spin much faster than they rotate around each other, e.g. $\omega_{p}\gg\zeta$. As we will see later, this assumption is reasonable for assemblies of magnetic colloids. Then the angular momentum remains a constant given by Eq. (10), while the vorticity is given by the equation:
$\nabla^{2}\zeta-\delta^{-2}\zeta=0,\quad\delta=\sqrt{\frac{\eta+\eta_{R}}{\Gamma_{v}}}.$ (11)
Appendix B Topological theory and edge solution
Notably, the operator in Eq. (11) has the same form as the square of a well-known topological Hamiltonian, the Dirac Hamiltonian, i.e., $(-\nabla^{2}+\delta^{-2})I_{2}=\mathcal{H}_{d}^{2}$, where $I_{2}$ is a rank two identity matrix [32]. The Dirac Hamiltonian is $\mathcal{H}_{d}=m\sigma_{1}-i\partial_{x}\sigma_{2}-i\partial_{y}\sigma_{3}$, where $\sigma_{1,2,3}$ are the Pauli matrices and $m=\pm\delta^{-1}$. The eigenvectors of $\mathcal{H}_{d}$ have a winding number known as the Chern number, which yields edge states robust to defects and random perturbations [23, 50].
At the interface between particle-rich and particle-poor regions, the Chern number changes its value, inducing a state with zero eigenvalue localized at the interface [32]. This state is exactly the steady state solution of Eq. (11), since the two operators $\mathcal{H}_{d}$ and $\mathcal{H}_{d}^{2}$ have the same zero eigenvalue solution. This solution yields a topologically-protected steady state where vorticity $\zeta(x)$ decays exponentially away from the edge, $x$.
Figure 5: (a) Colloidal cluster forming a semi-infinite plane. (b) Colloidal cluster with a generic boundary. At each point of the boundary define its curvature and a polar coordinate system.
To analytically find this exponentially decaying solution of Eq. (11), we adopt another simplification, namely that the boundary of colloidal assembly is close to being flat. Then we can consider the colloidal assembly as a semi-infinite plane at $x>0$ with $x=0$ line being its boundary (see Fig. 5(a)). Then, the solution is $y$-independent and is
$\zeta(x)=\zeta_{0}e^{-x/\delta}.$ (12)
@@ Appendix B Topological theory and edge solution
$\sigma_{yy}\big|_{x=0}$ $=-p=0,$ (13)
$\sigma_{xy}\big|_{x=0}$ $=2\omega_{p}\eta_{R}+(\eta+\eta_{R})\zeta_{0},$ (14)
@@
$\zeta_{0}=\frac{2\eta_{R}\omega_{p}}{\eta+\eta_{R}}.$ (15)
@@
$v_{y}(x)=-\int\limits_{x}^{\infty}\zeta(x)dx+v_{y}(\infty)=\frac{2\eta_{R}\omega_{p}\delta}{\eta+\eta_{R}}e^{-x/\delta}+v_{\mathrm{th}},$ (16)
@@
Figure 6: Edge flow profile modeled in a cluster of radius $R$ as a function of the radial coordinate $r$. Red line shows an exponential decay with decay length $\delta$. Blue dashed line shows a modified Bessel function of the first kind with decay length $\delta$. Both functions are normalized to their values at the edge of the cluster. The ratio $R/\delta=5$.
Appendix C Solution at curved boundary
Figure 7: Plot fitted angular velocity of the cluster bulk as a function of the cluster radius for three different magnetic field strengths. Solid lines show power law fits.
Figure 8: Left: mean speeds in a cluster at $H_{0}=1.1$ mT (black dots) are fitted by two models, without bulk rotation (red line) and with bulk rotation (purple line). Shaded areas indicate the 95% confidence intervals. Right: residuals are plotted for the fit without bulk rotation (top panel) and with bulk rotation (bottom panel). Including bulk rotation makes residuals more symmetric around zero and the sum of their squares smaller.
In this Appendix, we find the edge flow solution in a more general case of a curved boundary with a curvature $\kappa=1/R$. We introduce polar coordinates $r,\varphi$ with the center of coordinate located at the center of the circle with radius $R$ [cf. Fig. 5(b)]. We redefine linear velocity in these coordinates as $v_{r}=\dot{r}$ and $v_{\varphi}=r\dot{\varphi}$ and the vorticity is expressed as $\zeta=1/r\,v_{\varphi}+\partial_{r}v_{\varphi}-1/r\partial_{\varphi}v_{r}$. Rewriting Eq. (11) in polar coordinates we get
$r^{2}\partial_{r}^{2}v_{\varphi}+r\partial_{r}v_{\varphi}-(1+r^{2}/\delta^{2})v_{\varphi}=0,$ (17)
@@ Appendix C Solution at curved boundary
$v_{\varphi}(r)=\begin{cases}AI_{1}(r/\delta),&\textrm{convex boundary}\\ BK_{1}(r/\delta),&\textrm{concave boundary}\end{cases}$ (18)
@@
$\sigma_{rr}\big|_{r=R}=-p=\gamma/R,$ (19)
$\sigma_{\varphi r}\big|_{r=R}=$
$=(\eta_{R}+\eta)\partial_{r}v_{\varphi}+(\eta_{R}-\eta)/r\,v_{\varphi}-2\eta_{R}\omega_{p}\big|_{r=R}=0,$ (20)
@@
$v_{\varphi}(r)=\begin{cases}\frac{\eta_{R}I_{1}(r/\delta)}{\eta I_{2}(R/\delta)+\eta_{R}I_{0}(R/\delta)}2\omega_{p}\delta,&\textrm{convex}\\ -\frac{\eta_{R}K_{1}(r/\delta)}{\eta K_{2}(R/\delta)+\eta_{R}K_{0}(R/\delta)}2\omega_{p}\delta,&\textrm{concave}\end{cases}$ (21)
Note, that the modified Bessel function also forms an exponential decay. If the cluster radius is a few times larger than the decay length of the edge flow, the Bessel function is very close to the exponential decay, as we show in Fig. 6. All clusters and voids in our experiments have a radius to decay length ratio $R/\delta\gtrsim 5$, meaning they are large enough to be modeled by the exponential function in Eq. (2) in the main text.
Appendix D Derivation of bulk cluster rotation
In this Appendix, we derive the bulk angular velocity of a cluster induced by shear stress from the edge flows. We assume that only a thin outer layer of particles moves with velocity $v_{\varphi}(R)$. The friction between this layer and the bulk of the cluster imposes a torque
$\tau_{1}=2\pi R\cdot v_{\varphi}(R)\eta\cdot R,$ (22)